OpenStudy (anonymous):

Given the perimeter and side lengths, write and equation to represent the perimeter of the rectangle. Solve for x,showing your work. P=38 units. 3x-9. X

OpenStudy (triciaal):

perimeter is all the way around 2(length + width) 38 = 2{(3x-9) + x } x=7

OpenStudy (anonymous):

Can you show your work

OpenStudy (triciaal):

38 = 2(4X - 9) 38 = 8X -18 38 + 18 = 8X -18 + 18 56 = 8X 56/8 = 8X/8 X = 7 3X - 9 =3(7) -9 = 21-9 = 12 TO CHECK 12 +12 +7 +7 = 38